OpenStudy (anonymous):

How can you find out whether or not 8pi/9 lies within the interval 0

OpenStudy (michele_laino):

please note that 8/9 = 0.88 which is less that 1

OpenStudy (michele_laino):

or alternatively, we have: 8/9 = 1 - 1/9

OpenStudy (anonymous):

I was making it harder than that lol, thank you!

OpenStudy (michele_laino):

thank you!